# Sweep Angle Effects of Flow Over an Undulated Cylinder

### v1

@@ 3.3 Vortex Wake Effects
Figure 8: Vorticity magnitude averaged over the entire span in $\bm{z}$ and a vertical range of $\bm{y/T=0-2}$.
Figure  8 plots spatially averaged vorticity magnitude in the wake to enable a quantitative comparison of vortex structures. As ellipse geometry did not exhibit spanwise variation in wake structures with the exception of $\mathit{\Lambda}=60$, a single red line on each plot demonstrates the primarily $\mathit{z}$-oriented vorticity resulting from roller-type vortices. The whisker geometry, particularly for $\mathit{\Lambda}=0$, instead exhibits strong spanwise variation in shed vortices. Its span average is represented by a black curve and an additional shaded region displays one standard deviation above and below the span-averaged curve highlighting the large spanwise vorticity variation.
Comparing the vorticity magnitude plots to their isosurface counterparts, at $\mathit{\Lambda}=0$ roller vortices shed by an ellipse result in a periodic pattern of peaks in vorticity magnitude corresponding to isosurfaces of negative $\omega_{z}$. The vorticity magnitude of the whisker geometry departs significantly from this pattern in both the near and far wake. The disruption of rollers into the spanwise phase-shifted horseshoe vortex pattern results in a lower average vorticity magnitude close to the trailing edge than an ellipse and a misalignment of peaks with those observed for an ellipses’ roller vortices.
At $\mathit{\Lambda}=15$ the mean vorticity magnitude between the two geometries converges to a similar pattern of periodic peaks with the standard deviation decreasing compared to $\mathit{\Lambda}=0$. While some variation in peak heights remain, the ellipse span-average curve is generally contained within one standard deviation of the whiskers’ span-averaged magnitude. The gradual increase in wake similarity continues for $\mathit{\Lambda}=30-45$ as the increasing spanwise component of velocity causes the whisker geometry to more closely emulate the span-symmetric flow over an ellipse. Finally, at $\mathit{\Lambda}=60$ the transition in shedding patterns experienced by the ellipse results in a constantly decreasing mean vorticity magnitude after the near wake region. In contrast, the whisker geometry continues to demonstrate peaks consistent with its roller-type vortex street shedding pattern, features which will influence the distribution and dissipation of energy in the wake in later analyses and explain the difference in RMS lift coefficient at $\mathit{\Lambda}=60$ where the undulated cylinder exceeded the ellipse.
(a) $\mathit{\Lambda}=0,Q=0.6$
(b) $\mathit{\Lambda}=15,Q=0.6$
(c) $\mathit{\Lambda}=30,Q=0.4$
(d) $\mathit{\Lambda}=45,Q=0.32$
(e) $\mathit{\Lambda}=60,Q=0.16$
Figure 9: Isosurfaces of $\bm{Q}$-criterion colored by the sign of $Z$-vorticity for $Re$ = 500.
To enable a comparison, $Q$-criterion at $Re=500$ are shown in Figure 9. At this higher Reynolds number, all wakes demonstrate increased vortex strength and, for ease of of visualization, are plotted using larger $Q$ values than the $Re=250$ data. The wake also exhibits increased local turbulence demonstrated by the finer and more complex $Q$ structures. Additionally, spanwise instabilities characteristic of the flow regime are prevalent, particularly visible for the ellipse in the form of many streamwise finger-like structures along the span, and not present at $Re=250$. However, with these developments, the ellipse still maintains a vortex street shedding pattern up until angle $\mathit{\Lambda}$ = 60 where the larger component of spanwise velocity again transitions to a separate shedding regime as seen for $Re=250$.
Next, comparing a smooth ellipse and whisker geometry at this higher Reynolds number reveals additional explanations for the force results. First examining a smooth ellipse, particularly for angles $\mathit{\Lambda}$ = 0 and 15, the smooth geometry produces noticeably shortened visible wakes, forming stronger and more compact vortices close to the trailing edge surface reflecting the significantly increased RMS lift coefficient values observed for $Re=500$. In contrast, the undulated cylinder sheds vortices much further from the model surface and continues to exhibit a notable phase difference in the shedding of horseshoe-type vortex structures from adjacent wavelengths at ${\Lambda}=0$, while experiencing only a small increase in lift coefficient values during the flow regime change. Beyond this point, larger angle orientations once again exhibit increasing spanwise symmetry similar to observations of the $Re$ = 250 simulations as the whisker geometry RMS lift coefficient continually approaches the ellipse. Finally, in the case of $\mathit{\Lambda}$ = 60, the whisker geometry again appears to delay the transition of flow to a peeling-off regime, instead maintaining more roller-like structures which are reflected in the larger RMS lift coefficient compared to an ellipse at this orientation.
It is worth noting that other studies, e.g. [40], have extended the use of the $Q$-criterion by incorporating both positive and negative $Q$ values in combination with force partitioning methods to enable a deeper dissection of the interplay between vortex dynamics and force generation. These methods highlight the potential for future work in the area of whisker-inspired geometry, particularly in relation to the forces observed as a result three dimensional wake of this complex geometry.
@@
Figure 10: Contours of turbulent kinetic energy in the wake accompanied by plots of span-averaged TKE.
To further explore the effect of the modified flow patterns created by the whisker geometry, turbulent kinetic energy in the wake can be correlated with observed vortex structures. Figure 10 displays contours of TKE with respect to sweep angle in the $xz$-plane for $Re=250$. As observed in prior seal whisker literature [34], at $\mathit{\Lambda}=0$ the largest values of TKE occur in periodic spanwise regions that align with the streamwise vortex structures visible in Figure 7 and, compared to a smooth ellipse, exhibit a lower peak magnitude and form further downstream in the wake. At $\mathit{\Lambda}=15$, the influence of altered streamwise vortex structures is still visible in contours of TKE where their interference with adjacent structures creates long pairs of horizontal mixing regions.
Also of note is the significant increase in TKE magnitude for the entire sampled region which correlates with the increase in unsteady lift forces observed with the initial addition of sweep angle, contrasting the drop in forces and TKE seen for smooth geometry. Figure 10, column three, demonstrates this observation through span-averaged TKE as a function of $x/T$. In Figure 10(a), the curves representing the whisker and ellipse show a large decrease in TKE magnitude for the whisker geometry to levels below even that observed for $\mathit{\Lambda}=30$. TKE magnitude decreases with an addition of sweep angle at $\mathit{\Lambda}=15$ for the ellipse but initially jumps for the whisker geometry at the same angle, mirroring the behavior of $Re=250$ forces discussed previously. Beyond this point, the magnitude of wake TKE decreases steadily with sweep angle to minimum values at $\mathit{\Lambda}=60$.
The $x$ location at which these peak magnitudes occur is also strongly influenced by geometry prior to sweep and then becomes relatively uniform with increasing angle. Figure 10(a) demonstrates this increased formation length as the strongest contour levels form much closer to the model surface in the wake of a smooth ellipse. This is also reflected in the TKE magnitude plot where the $\mathit{\Lambda}=0$ whisker peak value occurs much further into the wake than the corresponding ellipse. With added sweep angle, this distance collapses quickly for the whisker as shown by Figure 10(b) where the highest contours of TKE appear at a similar $x$ location for both geometries and in the magnitude curves where the pea whisker TKE value occurs at a lower $x/T$ value that is closer to that of the ellipse peak.
For $\mathit{\Lambda}=30$ and beyond, the influence of undulations on increased streamwise vorticity and its disruptive effects on the formation of roller vortices diminishes. A periodic wake TKE pattern is still visible for all large sweep angles, however the magnitude with respect to spanwise location becomes increasingly uniform. For $\mathit{\Lambda}=45$, spanwise momentum transport is now dominated by spanwise flow instead of surface undulations and the resulting wake increasingly mirrors that of an ellipse with roller vortices and the accompanying pattern in TKE as they dissipate. At $\mathit{\Lambda}=60$, as was shown by flow structures in Figure 7, the presence of undulations appears to delay the transition of vortex shedding to a shallow angle regime with a noticeable mismatch from the sweep angle that is also reflected in contours of TKE.
@@ 4 Conclusion
Numerical simulations of flow over a whisker-inspired geometry are investigated at Reynolds numbers of 250 and 500 over a range of sweep angles. Comparing the whisker geometry to a smooth cylinder results in a significant in unsteady lift and drag forces across the entire range of tested sweep angles. To isolate the effects of the undulated surface, comparison with a smooth ellipse of equivalent dimensions ,which also experiences significantly lower forces than a a circular cylinder, reveals a reduction of both forces for the range of flow angles of $\mathit{\Lambda}$ = 0 to 45 while $\mathit{\Lambda}=60$ showed a slight increase in RMS lift over an ellipse. The impact of surface undulations is strongest in flow perpendicular to the model span with a powerful reduction in RMS lift force by 91.1% and 84.2% from that of a smooth ellipse for the lower and higher Reynolds numbers, respectively. In terms of drag forces, there is a reduction of 10.4% and 8.6%, respectively. As sweep angle increases forces are still reduced, albeit by a smaller magnitude, up to a sweep angle of $\mathit{\Lambda}=45$ where the ellipse and whisker converge to similar values before ultimately diverging with a larger whisker lift force at $\mathit{\Lambda}=60$.
The response of the whisker-inspired shape is influenced by the three-dimensional flow driven by its spanwise-varied cross-section demonstrated to have a complex evolution with respect to increasing angle. With larger sweep angle, as the whisker undulations increasingly approximate the relative spanwise symmetry in aspect ratio of an ellipse, the strong spanwise flow effects generated by surface undulations are observed to gradually give way to ellipse-like shedding behavior. This also mirrors the ellipse trend in forces and results in increasingly similar flow structures. While perpendicular flow demonstrated distinct shedding regions along the span visualized by a unique separation line and flow structures with strong streamwise vorticity, large sweep angles merged these shedding regions and approximated the flow structures of a smooth ellipse. The exception is that of the largest sweep angle tested at $\mathit{\Lambda}$ = 60. Here undulations delayed the transition to a different wake structure. These altered flow patterns were also reflected in contours of turbulent kinetic energy in the wake, which began with concentrated mixing regions that eventually spread and merge across the span similarly to an ellipse.
Based on resulting force measurements, angled whisker-inspired cylinder geometry for sweep angles of $\mathit{\Lambda}=30$ or smaller shows the most promise in applications of oscillating lift force and vortex induced vibration reduction and to a lesser extent, the reduction of drag force. Observations made at two bioinspired Reynolds number values indicate that the forces, shedding, and wake structure effects are not strictly sensitive to a specific flow condition within this regime. The extension of this geometry and its beneficial qualities to higher Reynolds numbers is still an open question.
@@ Appendix A Mesh Independence
The numerical grid used for these simulations consists of 4.14 million cells and is the same as as the mesh developed and validated in previous literature [5, 13, 34] for simulations at $Re$ = 500 after comparison to structured meshes containing 1, 2, and 6 million cells. Figure A1 shows that,, while still well-resolved, the addition of sweep to the domain produces a larger spanwise component of flow for which the original mesh was not validated.
@@
$N_{total}$ $N_{z}$ $N_{\theta}$ $N_{r}$ $\Delta z/T$ $\Delta r/T_{min}$ $\overline{C_{D}}$ $St$

### v2

@@ 3.3 Vortex Wake Effects
Figure 10: Contours of $\bm{Re=250}$ TKE in the wake accompanied by plots of span-averaged TKE.
Turbulent kinetic energy in the wake can be correlated with observed vortex structures, as shown by Figure 10 for $Re=250$. As previously observed at $\mathit{\Lambda}=0$ [lyons_wavelength_2022], the largest values of TKE occur in periodic spanwise regions that align with the streamwise vortex structures visible in Figure 8 and, compared to a smooth ellipse, exhibit a lower peak magnitude and form further downstream in the wake. At $\mathit{\Lambda}=15$, the influence of altered streamwise vortex structures is still visible in contours of TKE where their interference with adjacent structures creates long pairs of streamwise mixing regions.
There is also a strong increase in TKE magnitude, correlating the increase in unsteady lift, when sweep is introduced at 15 degrees. Figure 10, column three, demonstrates this observation through span-averaged TKE as a function of $x/T$. Figure 10(a) shows a large decrease in TKE magnitude for the whisker geometry. As sweep is introduced TKE magnitude decreases for the ellipse but initially jumps for the whisker geometry, mirroring the behavior of the force coefficients. Beyond this point, the magnitude of wake TKE decreases steadily with sweep angle to minimum values at $\mathit{\Lambda}=60$. The downstream $x$ location at which TKE peak magnitudes occur is also strongly influenced by geometry prior to sweep and then becomes relatively uniform with increasing angle.
For $\mathit{\Lambda}=30$ and beyond, a periodic TKE pattern is visible in the wake however the magnitude with respect to spanwise location becomes increasingly uniform. For $\mathit{\Lambda}=45$, spanwise momentum transport is now dominated by the freestream spanwise component instead of surface undulations, and the resulting wake increasingly mirrors that of an ellipse. At $\mathit{\Lambda}=60$, as is shown by flow structures in Figure 8, the presence of undulations appears to delay the transition of vortex shedding to a shallow angle regime with a noticeable mismatch from the sweep angle that is also reflected in contours of TKE.
At $Re=500$, shown in Figure 11, flow is more unstable and prone to three-dimensional breakdown, yet the overall spatial reorganization of TKE remains consistent with trends observed at $Re=250$. A smooth ellipse demonstrates an anticipated increase in peak intensity across all contours. The seal whisker demonstrates a similar periodic spanwise pattern at $\mathit{\Lambda=0}$ with a concentration of TKE in paired structures in the wake of each wavelength. With added sweep, this pattern is progressively disrupted, though periodic influence from the surface undulations remains evident. While sweep continues to redistribute TKE in the wake and reduce spanwise variation, an important Reynolds-number effect emerges. Span-averaged TKE shows that the difference between whisker and ellipse geometries not only persists, but becomes significantly larger than at $Re=250$, particularly for $\mathit{\Lambda\geq 30}$. In this high-sweep regime, force suppression was observed to diminish, yet the whisker maintains substantially lower wake TKE than the ellipse. This indicates that at higher Reynolds number, surface undulations become increasingly effective at redistributing turbulent energy into more three-dimensional convecting structures. Thus, even where force behavior converges, the whisker continues to alter wake turbulence intensity and organization relative to an ellipse.
@@
Figure 11: Contours of $\bm{Re=500}$ TKE in the wake accompanied by plots of span-averaged TKE.
3.4 Shedding Frequency Trends
(a) $Re=250$
(b) $Re=500$
Figure 12: Spectrum of unsteady lift forcing for all geometries and sweep angles at Reynolds number 250 and 500.
The differences in the shedding pattern shown by flow structures have the strongest influence on the shedding frequency spectrum of each geometry at smaller sweep. Spectra for the cylinder, ellipse, and whisker across all sweep angles are shown in Figure 12. As noted in prior work [dunt_frequency_2024], the $\mathit{\Lambda}=0$ whisker exhibits spanwise phase-shifted shedding with lower spectral magnitude and reduced peak frequency relative to smooth geometries. With the addition of sweep, a coherent shedding peak reemerges and the whisker shedding frequency converges toward the cylinder and ellipse values at $\mathit{\Lambda}=15$ and $30$, though amplitudes remain lower, consistent with reduced RMS lift. At $\mathit{\Lambda}=45$, the whisker spectrum surpasses the ellipse, matching the previously identified break-even point in RMS lift while arising from distinct wake dynamics. At this large sweep angle, the force-suppression capability of the whisker is diminished, and the dominant shedding frequency across geometries appears to be governed primarily by the effective bluff-body dimensions which are similar for an ellipse and whisker but distinct from the cylinder. Then, at $\mathit{\Lambda}=60$, the whisker is observed to delay the ellipse’s transition to a new shedding regime, retaining stronger periodic lift fluctuations while the ellipse exhibits almost entirely diminished oscillations. Overall, sweep angle causes a gradual realignment of whisker shedding and ellipse-like behavior before later demonstrating the key divergence between the two geometries for the largest sweep angle.
At $Re=500$, the shedding spectra show greater noise form the more transitioned wake, but key distinctions remain. The difference between peak frequencies across all geometries is initially greater, but all once again converge to similar values at $\mathit{\Lambda}=45$. Notably, the largest sweep angle $\mathit{\Lambda}=60$ demonstrates smoother spectra for all geometries representing a delayed introduction of three-dimensional flow effects and turbulent transition. Additionally, the smooth ellipse at this angle now exhibits a coherent peak as opposed to its diminished periodic shedding behavior at lower Reynolds number.
@@ 4 Conclusion
Numerical simulations of swept flow over a whisker-inspired geometry are investigated at Reynolds numbers of 250 and 500. A comparison between the whisker geometry and a smooth circular cylinder reveals a significant reduction in unsteady lift and drag forces across the entire range of tested sweep angles. Comparison with a smooth ellipse of equivalent aspect ratio reveals a reduction of forces for the range of flow angles of $\mathit{\Lambda}$ = 0 to 45 and a slight increase in RMS lift at $\mathit{\Lambda}=60$. The impact of surface undulations is strongest in flow perpendicular to the model span ($\mathit{\Lambda}=0$) with a reduction in RMS lift force by 86.1% and 90.8% from that of a smooth ellipse for $Re=250$ and $Re=500$, respectively. In terms of drag forces, there is a reduction of 9.0% and 11.4% as sweep angle increases forces are reduced, by a smaller magnitude, up to a sweep angle of $\mathit{\Lambda}=45$ where the ellipse and whisker converge to similar values before ultimately diverging with a larger whisker lift force at $\mathit{\Lambda}=60$.
The forces on the whisker are influenced by the three-dimensional flow caused by the spanwise-varied cross-section. This has a complex evolution with respect to increasing sweep angle since the relative cross-section is a function of $\mathit{\Lambda}$. With larger sweep angle, the whisker undulations increasingly approximate the relative aspect ratio of an ellipse, and thus the strong spanwise flow effects generated by surface undulations are observed to gradually give way to ellipse-like shedding behavior.
While perpendicular flow has distinct shedding regions along the span, large sweep angles merge these shedding regions and approximate the flow structures of a smooth ellipse. The exception is that of the largest sweep angle tested at $\mathit{\Lambda}$ = 60, where undulations delayed the transition to a different wake structure.
The results indicate that undulated cylinders with sweep angles of $\mathit{\Lambda}=30$ or smaller can effectively reduce RMS lift and the resulting vortex-induced vibration reduction, and to a lesser extent, can reduce drag force. Observations made at two bioinspired Reynolds number values demonstrated that sweep-induced force suppression and wake modifications are not strongly dependent on Reynolds number within this range. While higher Reynolds number flows exhibit increased unsteadiness and less coherent wake structures, the primary effects of surface undulations, suppression of vortex-induced RMS lift and convergence towards ellipse-like behavior with increasing sweep angle, persisted across both $Re$ values. This demonstrates that the hydrodynamic benefits of the whisker-inspired geometry are robust to realistic changes in swimming speed. However, analysis of span-averaged turbulent kinetic energy reveals that the relative wake energy reduction provided by the whisker geometry becomes more pronounced at $Re=500$, particularly at the largest sweep angles. This indicates that although force suppression diminishes with sweep, the whisker continues to alter wake structure in a manner that limits the persistence of energetic turbulent structures more effectively than an ellipse. Thus, wake stabilization effects of the undulated geometry extend beyond the range of effective force suppression which became comparable to an ellipse at a sweep angle of 45 degrees. This convergence reflects the balance between sweep-induced attenuation of forces and the diminishing influence of whisker-induced suppression, rather than a sharply defined biological threshold. In a real whisker array, which spans a continuum of effective orientations, this crossover would be distributed across whiskers and along their length. However, whiskers are more likely to occupy the range of orientations associated with stronger lift suppression when protracted forward into the flow, suggesting this behavior may enhance the effectiveness of their sensing abilities.
While the present study does not model whisker motion, upstream wakes, or fluid-structure interactions, the observed RMS lift trends suggest likely implications for VIV and hydrodynamic sensing. At small sweep angles, undulations strongly disrupt coherent vortex shedding, reducing unsteady forces and indicating the greatest benefit for vibration suppression. As sweep increases, undulation effects diminish and the wake becomes more coherent, but the absolute magnitude of unsteady forces is lower, so background forcing remains moderate. These results imply that whiskers achieve the greatest VIV reduction and potentially enhanced sensing when protracted forward. The persistence of suppression across a broader range of orientations highlights functional robustness to natural variation in whisker angle and environmental flow. By reducing background unsteady forces, whiskers remain relatively still in mean flow, which would allow incoming disturbances from upstream wakes or objects to produce more noticeable perturbations, consistent with theories that this mechanism underlies their effectiveness as hydrodynamic sensors and that this effect persists with the addition of sweep angle.
These findings highlight the importance of sweep angle as a biologically-informed and physically significant parameter. Quantitative assessment of VIV and sensing performance will require further investigation incorporating flexible whiskers or interactions with upstream objects, and the extension of this geometry and its beneficial qualities to higher Reynolds numbers remains an open question.
@@ 5 Acknowledgments
We thank Dr. Colleen Reichmuth, the Pinniped Cognition and Sensory Systems Laboratory and Thia Elliott for imagery of the harbor seal used in Figure 1, and Sprouts the seal for the insight and inspiration that he provided to this research.
@@ Appendix A Mesh Independence
The numerical grid used for these simulations consists of 4.14 million cells and is the same as the mesh developed and validated in previous literature [lyons_flow_2020, yuasa_simulations_2021, lyons_wavelength_2022] for simulations at $Re$ = 500 after comparison to structured meshes containing 1, 2, and 6 million cells. Figure A1 shows that, while still well-resolved, the addition of sweep to the domain produces a larger spanwise component of flow for which the original mesh was not validated.
@@
$N_{total}$ $N_{z}$ $N_{\theta}$ $N_{r}$ $\Delta z/T$ $\Delta r/T_{min}$ $\overline{C_{D}}$ StSt